OpenStudy (help!!!!):

The domain of the function y(x) =3 cos(5x-4) +1 is all real numbers. Which of the following is the range of the function y(x) ?

OpenStudy (help!!!!):

Here are the answer choices \[a. -3\le y(x) \le 3\] \[b. -4\le y(x) \le 3\] \[c. -4 \le y(x) \le 2\] \[d. -2 \le y(x) \le 4\]

OpenStudy (help!!!!):

The answer is D but I need an explanation..

OpenStudy (help!!!!):

@mathstudent55 @zepdrix @dan815 @Redcan @tom982 @whpalmer4 @inkyvoyd @pooja195

OpenStudy (sparrow2):

you just need to keep in mind that cosx takes value form -1 to 1

OpenStudy (sparrow2):

so for least value it is -1 and if you insert you get : 3*(-1)+1=-2 for the maximum it is 1 and you have 3*1+1=4 so [-2;4]

OpenStudy (help!!!!):

@sparrow2 how did u know that it was cosx? Sorry I new to this kind of material

OpenStudy (sparrow2):

from your equation it is cos(5x-4) it...argument doesn't matter the range of cosx is [-1,1]

OpenStudy (victoriasushchik):

cos(t) will be, at most 1 and at least -1 3 * (-1) + 1 = -3 + 1 = -2 3 * (1) + 1 = 3 + 1 = 4 So the range is [-2 , 4]

OpenStudy (help!!!!):

Thanks for the clarification!

OpenStudy (victoriasushchik):

np =] medal + fan me thankz